OpenStudy (aravindg):

qn below:

OpenStudy (aravindg):

\[\large v_n=\dfrac{e}{\sqrt{4 \pi \in mr_n}}\] \[\large m v_n r_n=\dfrac{nh}{2 \pi}\] I need to get expression for v_n and r_n combining these equations .

OpenStudy (sirm3d):

divide the second by the first. you should get an equation in r_n only

OpenStudy (koikkara):

@sirm3d ...y not solve in "draw" and help him....

OpenStudy (sirm3d):

why don't you draw it, @Koikkara ? feel free to join us.

OpenStudy (aravindg):

but the answer i have doesnt match \[r_n=(\dfrac{n^2}{m})(\dfrac{h}{2 \pi }^2 )\dfrac{4 \pi \in }{e^2}\]

OpenStudy (aravindg):

\[v_n=\dfrac{1}{n} \dfrac{e^2}{4 \pi \in} \dfrac{1}{(h/2 \pi)}\]

OpenStudy (sirm3d):

\[\Large{\frac{m^2v_n^2r_n^2}{v_n^2}=\frac{n^2h^24\pi\varepsilon m r_n}{4\pi^2e^2}\\r_n=\frac{n^2h^2\varepsilon}{e^2m}=\frac{n^2}{m}\frac{h^2}{\pi}\frac{\varepsilon }{e^2}}\]

OpenStudy (sirm3d):

if you multiply the middle fraction by 4pi/4pi, \[\Large \frac{h^2}{\pi}=\frac{h^24\pi}{4\pi^2}=\left(\frac{h}{2\pi}\right)^24\pi\]

OpenStudy (aravindg):

oh i seee..then what about v_n ?

OpenStudy (sirm3d):

in the second equation \[\Large v_n=\frac{nh}{2\pi m}\cdot \frac{1}{r_n}\]

OpenStudy (aravindg):

then how i bring the other terms ?

OpenStudy (aravindg):

should i substitute for r_n ?

OpenStudy (sirm3d):

\[\Large v_n=\frac{nh}{2\pi m}\frac{m}{n^2}\left(\frac{2\pi}{h}\right)^2\frac{e^2}{4\pi\varepsilon}=\frac{e^2}{2nh\varepsilon }\]

OpenStudy (aravindg):

ok thanks a lot ..i spent a lot of time on this ....

OpenStudy (sirm3d):

you can rewrite the denominator as \[\Large 2nh\varepsilon =n(4\pi)(\frac{h}{2\pi})\varepsilon\]

OpenStudy (aravindg):

ok

OpenStudy (sirm3d):

this symbol \(\varepsilon\) is \varepsilon in TeX

OpenStudy (aravindg):

ok ... does this method of combining equations like these have a name ?

OpenStudy (aravindg):

i am coming across this type of solving for first time

OpenStudy (sirm3d):

just manipulation of symbols.

OpenStudy (sirm3d):

if algebra, if a=b and c=d, then a/c = b/d provided none of the denominators is zero.

OpenStudy (aravindg):

ok thanks !